Diffusion Morning Report 12/13/06

Ted Berg tberg at stanford.edu
Wed Dec 13 06:34:39 PST 2006


Good Morning All,
    Looking good this morning. LTO is down for scheduled pull and clean. 
Still looking for leaks in Epi.. AMAT process to be here today to 
continue processing.
POs for AMAT and TEL written this Morning.  No other problems at this time.

Diffusion 	  	  	  	  	 
  	Date: 	12/13/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	9.25/30 	  	 
Tylan 8-SiGe 	Up 	  	10.86/30 	  	 
Tylan 9-Poly 	Up 	  	19.38/30 	  	 
Tylan 10-Nitride 	Up 	  	28.65/30 	  	 
Tylan 11-BPSG-LTO 	Down 	Pull and Clean
	0/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Up 	  	  	  	 
Epi 	down 	leak testing 	  	  	 
  	
	
	
	
	
Burn Box 	up 	  	  	  	 
ZYGO 	Up 	  	  	  	 
  	up 	  	   	  	 
GAAS 22 	  	  	  	  	 
GAAS 23 	UP 	  	  	  	 
SiH4 for furnaces 	up 	  	  	  	 

	  	A on line 540 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20061213/69301c77/attachment.html>


More information about the diffusion mailing list